# Automatic Depression Assessment using Machine Learning: A Comprehensive Survey

### v1

@@ Abstract.
Automatic Depression Assessment, Machine Learning, Depression-related Human Behaviours
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXjournal: POMACSjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Computing methodologies Activity recognition and understandingccs: Applied computing Psychology
1. Introduction
@@
Depression, also called major depressive disorder (MDD), is a significant mental disorder that negatively impacts around 3.8% of the world population, i.e., with one in 15 adults diagnosed with depression every year (Thapar et al., 2022), which would physiologically, mentally and socially impair individuals, including damaging brain structures (van Eijndhoven et al., 2013; Lu et al., 2012) and perception ability (Bourke et al., 2010), causing chronic physiological illness (Gan et al., 2014), as well as substantial financial costs (Thapar et al., 2022; Lépine and Briley, 2011). Depression can be categorised into several types, including clinical depression (major depressive disorder, MDD), persistent depressive disorder (PDD), disruptive mood dysregulation disorder (DMDD), premenstrual dysphoric disorder (PMDD) and depressive disorder triggered by other medical conditions  [^1] https://my.clevelandclinic.org/health/diseases/9290-depression. The predominant characteristics of depression are the depressed mood or dysphoria reflected by various daily symptoms such as diminished interest in daily activities, alterations in sleep and appetite, feelings of guilt and hopelessness, fatigue, difficulties in concentration, and even suicide attempts (Abbey and Garfinkel, 1991; Paykel, 1977; Benca and Peterson, 2008).
@@
Existing clinical depression assessments typically rely on structured or semi-structured interviews (SDI) conducted by professional psychologists (Smith et al., 2013), where specifically-designed multi-point questionnaires (e.g., Hamilton Rating Scale for Depression (HAM-D) (Hamilton, 1960) and PRIME-MD Patient Health Questionnaire (PHQ) (Kroenke et al., 2001)) are required to be filled by either patients or psychologists to describe depression symptoms (Fletcher, 2008). However, these strategies frequently suffer from: (i) substantial costs charged by psychologists and clinical equipment usage; (ii) long assessment period caused by complex protocols (e.g. psychologist appointment scheduling and traveling); (iii) subjective assessment outcomes caused by disparities in psychologists’ expertise; and (iv) biased outcomes as patients may exaggerate/deemphasise crucial facts or symptoms in verbal reports/questionnaires. Such limitations can cause delays and unreliability in depression assessment.
@@
As a result, there is a recognised necessity within the medical community to explore cheap, fast, objective (repeatable) and reliable solutions for depression assessment, e.g., incorporating machine learning (ML) techniques for automatic depression assessment (ADA). These ADA approaches are generally built on consistent biological and physiological findings that various human internal and external behaviours are informative for reflecting depression status in clinical trails. Specifically, multiple brain behaviours (e.g., abnormal grey matter changes (van Eijndhoven et al., 2013)), physiological behaviours (e.g., ‘accelerated aging’ (Wolkowitz et al., 2010) and body shape changes (Milaneschi et al., 2019; Fu et al., 2023)) as well as external expressive behaviours such as facial behaviours (e.g., reduced expressions (Gupta, 2009)), gestures (Hinchliffe et al., 1975) (e.g., fewer gestures (Hinchliffe et al., 1975)), body movements (Deligianni et al., 2019; Gimeno-Gómez et al., 2024) verbal language behaviours (Nilsonne et al., 1988; Pennebaker et al., 2001) and non-verbal vocal behaviours (e.g., reduced pitch range (Hollien, 1980)), have been shown to indicate or be associated with depression.
@@
The majority of existing ADA solutions infer depression from easily accessible human external expressive behaviours (e.g., audio, facial and language behaviours), including traditional ML-based solutions (Kaya et al., 2014; Pampouchidou et al., 2016; Muzammel et al., 2021) that combine hand-crafted audio, visual or audio-visual features (e.g., Local Binary Pattern (LBP) and Mel-frequency cepstral coefficients (MFCCs)) with traditional ML regressors/classifiers (e.g., Support Vector Machine Regression (SVR), Random Forest (RF), Logistic Regression (LR) and k-nearest neighbor (KNN)), and DL-based ADA approaches (Jung et al., 2024; Xu et al., 2024; Niu et al., 2020; Khowaja et al., 2024; Ye et al., 2024; Fang et al., 2023a) leveraging Convolution Neural Networks (CNNs), Transformers or Graph Neural Networks (GNNs) for end-to-end depression feature extraction and assessment. To address ethical and privacy issues, in some occasions only de-identified behaviour data are provided, and thus some ADA approaches (Gahalawat et al., 2023; Abbasi et al., 2022; Pan et al., 2023; Gong and Poellabauer, 2017; Rodrigues Makiuchi et al., 2019; Wang et al., 2024a) were specifically developed on anonymous facial behaviour primitives (e.g., facial Action Units (AUs) and head poses) and de-identified audio/text signals. Among existing expressive behaviour-based ADA approaches, majority assess depression status based on an entire given clip (usually less than 30 minutes), while only a few studies (Girard et al., 2013; Ahmad et al., 2021; Shang et al., 2021) infer depression from static face images as long-term behaviours (typically lasting several weeks) may be more reliable indicators of depression (Song et al., 2022a; de Melo et al., 2021). Alternatively, other ADA solutions were developed for analysing brain behaviours recorded by non-invasive electroencephalography (EEG) (Zhang et al., 2024b; Shen et al., 2023) or magnetic resonance imaging (MRI) (Chen et al., 2021b; Gao et al., 2023; Mousavian et al., 2021). Similarly to external expressive behaviour-based ADA approaches, two types of pipelines have been explored: (i) hand-crafted pipeline that applies traditional ML predictors to process manually defined features (e.g., spectral EEG features reflecting depression-related human brain electrical activities (De Melo et al., 2019; Hosseinifard et al., 2013)); and (ii) DL-based pipeline that extends CNNs (Seal et al., 2021; Sharma et al., 2023; Gao et al., 2023; Chen et al., 2021b), Transformers (Shao et al., 2024b; Wang et al., 2024b) and GNNs (Lu et al., 2024; Li et al., 2023; Dai et al., 2023; Zhu et al., 2023) to identify depression-related dynamics and relationships within/among different EEG channels, or brain structural/behavioural abnormalities in MRI scans.
@@
In particular, a large part of existing expressive behaviour-based ADA approaches were developed on the AVEC depression challenge series (Valstar et al., 2013, 2014, 2016; Ringeval et al., 2017, 2019) organised since 2013. These challenges provided several mainstream publicly available audio-visual depression datasets facilitating researchers to share and compare their ADA solutions, including AVEC 2013 and AVEC 2014 audio-visual depression datasets containing human face videos and audio (in German language) files, as well as DAIC-WOZ dataset (Gratch et al., 2014) containing de-identified audio signals, facial primitives and language transcripts (used for AVEC 2016, AVEC 2017 and AVEC 2019 challenges). Other audio-visual datasets including CMDC (Zou et al., 2022), PCD (Dibeklioğlu et al., 2018), LaB (Lin et al., 2021), and D-vlog (Yoon et al., 2022) also have been made publicly available, where the LaB dataset additionally provides human body movements. All these datasets (except D-vlog (Yoon et al., 2022)) were recorded under pre-defined tasks or interviews, and provided rigorous clip-level depression labels. Besides, the EATD-Corpus (Shen et al., 2022) recorded audio and text transcript files while the MODMA dataset (Cai et al., 2022) recorded EEG and audio signals. Due to ethical considerations, to the best of our knowledge, most MRI/EEG and expressive behaviour depression datasets for ADA research are not publicly available.
@@
Given the numerous ML-based ADA approaches developed in the past decade, this paper conducts a comprehensive survey for up-to-date ML-based automatic depression assessment (ADA) approaches building on various internal human brain behaviours recorded by EEG and MRIs, as well as external human expressive speech, facial, body and language behaviours. While early ADA surveys (Gao et al., 2018; Pampouchidou et al., 2016; He et al., 2022b; Aleem et al., 2022) fail to include up-to-date advanced ADA approaches, recent surveys suffer from the following limitations: (i) reviewing only speech-based ADA solutions (Wu et al., 2023); (ii) reviewing audio and language-based solutions without providing comprehensive reviews for visual ADA approaches (Squires et al., 2023); and (iii) summing ADA approaches that analyse social media cues, ignoring audio, video, MRI and EEG-based ADA solutions (Squires et al., 2023). More importantly, none of these surveys provided comprehensive psychological/physiological/medical evidence linking depression to various human internal and external behaviours. Our survey starts with reviewing the biological, physiological and psychological evidence linking depression with a wide variety of human internal signals and external behavioural cues (Sec. 2.2), which provide a solid theoretical basis for developing human internal and external cue-based ADA approaches (surveyed in Sec. 4). In addition, publicly available ADA datasets are discussed in Sec. 5. Finally, Sec. 6 discusses the challenges that are faced by existing ADA approaches and highlights opportunities for future research. Fig. 1(a) illustrates statistics and trends of ADA publications in this survey, with detailed discussion provided in the Supplementary Material.
Figure 1. (a) Statistics and trend of ADA approaches (2012-2024) surveyed in this paper. (b) The negative impacts of depression.
2. Background
@@
2.1. Health and social impacts of depression
@@
As summarised by Fig. 1(b), depression affects both human health and their social functioning. (1) Health impacts: Depression frequently causes damages on human brain regions (Siegle et al., 2002), particularly in the hippocampus, cortical regions, frontallobes, and thalamic (Yang et al., 2023b). It also affects human immune systems (Eyre and Baune, 2012), leading to physiological illness such as coronary heart diseases (Gan et al., 2014) and cytokines-related illness (Yirmiya et al., 2000). A systematic survey (Katon et al., 2007) concludes that depressed chronically ill patients (e.g., diabetes, pulmonary disease and arthritis) suffer from significantly higher number of medical symptoms compared to non-depressed patients. Depression is also a high risk factor leading to peripheral vascular disease (d’Audiffret et al., 2010), which is also linked to chronic fatigue (Abbey and Garfinkel, 1991), insomnia (Benca and Peterson, 2008), atherosclerosis (Isingrini et al., 2011), Parkinson’s disease (Hemmerle et al., 2012) and cancers (andersen2010biobehavioural). (2) Societal impacts: Depression also leads to serious social problems. More than half of the suicide cases in Europe are linked to depression-related mental disorders (Hegerl, 2016), with another study (Dong et al., 2019) revealed similar findings. Depression also typically leads to poor interpersonal relationships (Stoetzer et al., 2009; Kupferberg and Hasler, 2023). The inadequate interpersonal relationships, in turn, signify diminished social support and contribute to suicide-related behaviours (Lewinsohn et al., 1997). Besides, addictive disorders are also frequently presented within depressed adolescents (Rao, 2006). Moreover, depressed individuals frequently suffer from substantial financial costs in the form of time away from work and a rise in healthcare expenses (Thapar et al., 2022), who are more susceptible to diminished earnings/job loss, homeless and increased substance abuse and poverty (Kupferberg and Hasler, 2023).
2.2. Depression-related human behaviours
2.2.1. Brain cues
@@
Depressed individuals typically experience abnormal brain processes. This is evidenced by brain regions implicated in the pathophysiology of depression (Mayberg et al., 2000). In particular, magnetic resonance imaging (MRI) have been widely employed to explore brain activities related to depression, revealing that depression would seriously impair several brain regions, including abnormal gray matter changes of dorsolateral prefrontal cortex (DLPFC) (Siegle et al., 2002), amygdala (Lu et al., 2012) and cerebellum (He et al., 2017), as well as damage symptoms of hippocampus and thalamus (significantly smaller volumes (van Eijndhoven et al., 2013; Yang et al., 2023b)). Some EEG studies found that depressed individuals usually exhibit hemispheric asymmetry in their brain signals compared to non-depressed individuals (Gotlib, 1998), as they usually have more alpha activity at the left side of their brains (Lee et al., 2018). Besides, Li et al. (Li et al., 2021) apply microstate analysis on EEG data of depressed and healthy individuals, suggesting that depression leads to increase in microstate C and decrease in microstate D, which further convinces depression’s negative influences on brain alpha band.
2.2.2. Physiological cues
Depression is also associated with abnormal physiological states, as changes in the brain affects on various body systems. For instance, fluctuations in serotonin that are related to depression frequently cause depressed individuals suffering from sleep problems (Benca and Peterson, 2008) and lower sex drive (Staley et al., 2006). The state of ‘accelerated aging’ (Wolkowitz et al., 2010) is another typical depression biomarker, as depression causes comparable changes on human brain’s plasticity and synaptic functions (Sibille, 2013) that are similar to those occurring with aging, where depressed individuals usually suffer from 4–6 years of accelerated aging compared to normal people (Verhoeven et al., 2014).
2.2.3. Physical indicators
Meanwhile, Katon et al. (Katon, 2003) and Lin et al. (Lin et al., 2011) concluded that depression leads individuals to experience decreased heart rate variability, increased adhesiveness of platelets, negative health habits such as smoking and over-eating, and low circadian rhythmicity in their skin body temperatures (Barbini et al., 1998). Appetite change is another physical indicator of depression, where greater appetite change tends to indicate greater depression severity (Paykel, 1977). As a result, human body shape and weight changes are important physical indicators to indicate depression (Polivy and Herman, 1976), e.g., the obesity (Milaneschi et al., 2019), where a reciprocal and reinforcing cycle of maladaptive physiological adjustments can establish an interconnection between the depression and obesity (Fu et al., 2023). As discussed in Sec. 2.1, sickness-like behaviours, cognitive dysfunction, and anhedonia (Eyre et al., 2013) are also typical depression-related cues. For example, there is a bidirectional relationships linking depression to the progression of cancer (Giese-Davis et al., 2011) and linking the persistent inflammatory processes to stresses in the manifestation of depressive symptoms (Sotelo et al., 2014).
2.2.4. Non-verbal expressive behaviours
Gesture and body movements: Depressed individuals typically show physical activity known as psychomotor retardation (Gupta, 2009), represented by different gait from normal people (Deligianni et al., 2019) such as slumped, hunched, or rigid posture that restricts body movements (France and Kramer, 2001)), impaired gesture movements (e.g., gesture deficits and degraded intransitive behaviours) (Pavlidou et al., 2021), fewer gestures during interpersonal interactions (Hinchliffe et al., 1975), as well as downwards head-tilting and frequent self-adaptor, self-touch or hand-over-face gestures (Mahmoud and Robinson, 2016, 2011; Lin et al., 2021).
Facial behaviours: A typical facial depression biomarker is the reduced or even diminished facial expressions (Gupta, 2009; Gaebel and Wölwer, 2004) and head movements (Joshi et al., 2013b), especially significantly reduced expressions of positive emotions. However, contradicting conclusions have been reached in terms of negative facial expressions, where some studies claimed that depressed individuals experience increased negative expressions (Ellgring, 2007) while others shown that even negative facial expressions are less expressed by them (Gaebel and Wölwer, 2004). From a more fine-grained perspective, Cohn et al. (Cohn et al., 2009) and Girard et al. (Girard et al., 2014) conducted pioneering investigations regarding the depression-related facial AU patterns, both of which show that participants suffering from high depression severity display less affiliative facial expressions (AU 12 and AU 15) but more non-affiliative facial expressions (AU 14). Also, an early study (Waxer, 1974) suggested that depressed individuals tend to avoid eye contact, and show more of nonspecific gaze and less mutual gaze, indicating lack of interest in communication.
Audio behaviours: Non-verbal audio behaviours are also crucial depression biomarkers (Tolkmitt and Scherer, 1986), as depression-related cognitive impairments would affect human working memory that controls the articulatory system and speech-based information storage (Murphy et al., 1999). Specifically, depressed individuals typically exhibit monotonous speech (Gupta, 2009) characterised by frequent pauses with a ‘muffled’ voice (Silva et al., 2021), reduction in articulatory coordination (Williamson et al., 2013) and verbal fluency (Husain et al., 2020), abnormal pitch variability, vocalization time, fundamental frequency and formants (Mundt et al., 2012; Nilsonne et al., 1988), long pause times (Mundt et al., 2012), as well as an overall decrease in speech rate/speed (Hollien, 1980). An early study summarised depressive voices as ‘patients speak in a low voice, slowly, hesitatingly, monotonously, sometimes stuttering, whispering, try several times before they bring out a word, become mute in the middle of a sentence’ (Kraepelin, 1921). Besides, Hollien et al. (Hollien, 1980) found that depressed individuals are more likely to experience reduced speaking intensity, pitch range and intonation, as well as a lack of linguistic stress, while Breznitz et al. (Breznitz, 1992) revealing that narrower frequencies are expressed by depressed females’ speeches compared to health females.
2.2.5. Verbal language behaviours
Language-based communication is a primary mean for people to exchange their ideas and intentions, and thus its styles are strongly associated with their depression status (Nilsonne et al., 1988), despite that there exists cross-cultural differences in depression-related verbal expressions (Tsugawa et al., 2015) (e.g., specific negative emotions or phrases are stronger depression indicators in Chinese social platforms than in other western platforms (Wang et al., 2013)). Among them, the most recognised verbal depression biomarkers are the increase usage of the first person singular (i.e., ‘I’, ‘me’, and ‘mine’) (Weintraub, 1989), swear words and depression terms, as well as the decrease usage of 3rd person pronoun (e.g., ‘She’, ‘He’). These are evidenced by the studies conducted on suicidal poets (Stirman and Pennebaker, 2001), student essays (Rude et al., 2004), Twitter (De Choudhury et al., 2013b) and a Chinese social platform - Sina (Wang et al., 2013), Other studies (De Choudhury et al., 2013a) further found that tentative words (e.g., maybe, perhaps), negations (e.g., no, never), adverbs, and conjunctions are more commonly used by depressive individuals, with less usage of verbs, while postpartum depression mothers tending to show more non-fluency in their social platform posts. From another perspective, Rude et al. (Rude et al., 2004) found that depressed individuals used more negatively valenced words, while Cummins et al. (Cummins et al., 2015; Gupta, 2009) summarised that they are more likely to express dull, misery and “lifeless” descriptors. Besides, the sentences expressed by depressed patients are usually fragmented or shortened, reflecting their reduced coherence and cognitive difficulty (Resnik et al., 2015).
2.2.6. Other depression-related human indicators
In addition to the behaviours discussed above, depressed patients also show various cognitive deficits compared to non-depressed ones, including varied exploratory behaviours (Blanco et al., 2013), impaired recognition of basic emotional facial expressions (Bourke et al., 2010), longer reaction time, a decreased psychomotor ability (Marazziti et al., 2010) and diminished capability in acquiring new information (but retaining the capability in gathering information) (Marazziti et al., 2010). Memory abnormality is a key sign reflecting depression, as depression causes the suppression of hippocampal neurogenesis, inhibition of dopamine neurons, and increased sensitivity of the amygdala (Dillon and Pizzagalli, 2018). As evidenced by (Ramponi et al., 2004), depressed individuals tend to exhibit difficulties in recalling information, i.e., they frequently experience enhanced memory for negative content but poorer memory for positive material (Burt et al., 1995). With such cognitive and memory deficits and bias, multiple human cognitive capabilities are further negatively impacted, i.e., depressed individuals frequently show abnormality or lower performances in decision-making (Blanco et al., 2013), problem-solving (Elderkin-Thompson et al., 2006), planning (Elliott et al., 1997) behaviours, as well as spending time with similarly depressed others and pair-wise interactions rather than group interactions (Elmer and Stadtfeld, 2020).

### v2

@@ Abstract
Index Terms:
Automatic depression assessment (ADA), Machine Learning, Deep Learning, Multi-modal analysis, Brain activities, Non-verbal behaviours, Language behaviours
1 Introduction
@@
Depression, also called major depressive disorder (MDD), is a significant mental disorder that negatively impacts around 3.8% of the world population, ranging from mild to severe conditions with hallucinations and delusions, i.e., with one in 15 adults diagnosed with depression every year [1], which would physiologically, mentally and socially impair individuals, including damaging brain structures [2, 3] and perception ability [4], causing chronic physiological illness [5], as well as substantial financial costs [1, 6]. Depression can be categorised into several types, including clinical depression (major depressive disorder, MDD), persistent depressive disorder (PDD), disruptive mood dysregulation disorder (DMDD), premenstrual dysphoric disorder (PMDD) and depressive disorder triggered by other medical conditions  [^1] https://my.clevelandclinic.org/health/diseases/9290-depression. The predominant characteristics of depression are the depressed mood or dysphoria reflected by various daily symptoms such as diminished interest in daily activities, alterations in sleep and appetite, feelings of guilt and hopelessness, fatigue, difficulties in concentration, and even suicide attempts [7, 8, 9].
@@
As a result, objective and accurate depression screening is crucial for raising early assessment and intervention. Existing clinical depression assessments typically rely on structured or semi-structured interviews conducted by professional psychologists [10], where specifically-designed multi-point questionnaires (e.g., Hamilton Rating Scale for Depression (HAM-D) [11] and PRIME-MD Patient Health Questionnaire (PHQ) [12]) are required to be filled by either patients or psychologists to describe depression symptoms [13]. However, these strategies frequently suffer from: (i) substantial costs charged by psychologists and clinical equipment usage; (ii) long assessment period caused by complex protocols (e.g. psychologist appointment and traveling); (iii) subjective assessment outcomes caused by disparities in psychologists’ expertise; and (iv) biased outcomes as patients may exaggerate/deemphasise crucial facts or symptoms in verbal reports/questionnaires. Such limitations can cause delays and unreliability in depression assessment, prolonging depression patients’ suffering.
@@
As a result, there is a recognised necessity within the medical community to explore cheap, fast, objective (repeatable) and reliable solutions for depression assessment, e.g., incorporating machine learning (ML) techniques for automatic depression assessment (ADA). These ADA approaches are generally built on consistent biological and physiological findings that various human internal and external behaviours are informative for reflecting depression status in clinical trails. Specifically, multiple brain behaviours such as abnormal grey matter changes [2], physiological behaviours (e.g., ‘accelerated aging’ [14] and body shape changes [15, 16]), as well as external expressive behaviours including facial behaviours (e.g., reduced expressions [17]), gestures (e.g., fewer gestures [18]), body movements [19, 20], verbal language behaviours (e.g., frequent first person singular usage [21, 22]), and non-verbal vocal behaviours (e.g., reduced pitch range [23]), can indicate or be associated with depression.
@@
The majority of existing ADA solutions infer depression from easily accessible human expressive behaviours (e.g., audio, facial and language behaviours), including traditional ML-based solutions [24, 25, 26] that combine hand-crafted audio, visual or audio-visual features (e.g., Local Binary Pattern (LBP) and Mel-frequency cepstral coefficients (MFCCs)) with traditional ML regressors/classifiers (e.g., Support Vector Machine Regression (SVR), Random Forest (RF), Logistic Regression (LR) and k-nearest neighbor (KNN)), and DL-based ADA approaches [27, 28, 29, 30, 31, 32] leveraging Convolution Neural Networks (CNNs), Transformers or Graph Neural Networks (GNNs) for end-to-end depression feature extraction and assessment. To address ethical and privacy issues, in some occasions only de-identified behaviour data are provided, and thus some ADA approaches [33, 34, 35, 36, 37, 38] were specifically developed on anonymous facial behaviour primitives (e.g., facial Action Units (AUs) and head poses) and de-identified audio/text signals. Among existing expressive behaviour-based ADA approaches, majority assess depression status based on an entire given clip (usually less than 30 minutes), while only a few studies [39, 40, 41] infer depression from static face images as long-term behaviours (typically lasting several weeks) may be more reliable indicators of depression [42, 43]. Alternatively, other ADA solutions were developed for analysing brain behaviours recorded by non-invasive electroencephalography (EEG) [44, 45] or magnetic resonance imaging (MRI) [46, 47, 48]. Similarly to external expressive behaviour-based ADA approaches, two types of pipelines have been explored: (i) hand-crafted pipeline that applies traditional ML predictors to process manually defined features (e.g., spectral EEG features reflecting depression-related human brain electrical activities [49, 50]); and (ii) DL-based pipeline that extends CNNs [51, 52, 47, 46], Transformers [53, 54] and GNNs [55, 56, 57, 58] to identify depression-related dynamics and relationships within/among different EEG channels, or brain structural/behavioural abnormalities in MRI scans.
@@
In particular, a large part of existing expressive behaviour-based ADA approaches were developed on the AVEC depression challenge series [59, 60, 61, 62, 63] organised since 2013. These challenges provided several mainstream publicly available audio-visual depression datasets facilitating researchers to share and compare their ADA solutions, including AVEC 2013 and AVEC 2014 audio-visual depression datasets containing human face videos and audio (in German language) files, as well as DAIC-WOZ dataset [64] containing de-identified audio signals, facial primitives and language transcripts (used for AVEC 2016, AVEC 2017 and AVEC 2019 challenges). Other audio-visual datasets including CMDC [65], PCD [66], LaB [67], and D-vlog [68] also have been made publicly available, where the LaB dataset additionally provides human body movements. All these datasets (except D-vlog [68]) were recorded under pre-defined tasks or interviews, and provided rigorous clip-level depression labels. Besides, the EATD-Corpus [69] recorded audio and text transcript files while the MODMA dataset [70] recorded EEG and audio signals. Due to ethical considerations, to the best of our knowledge, most MRI/EEG and expressive behaviour datasets for ADA research are not publicly available.
@@
Given the numerous ML-based ADA approaches developed in the past decade, this paper conducts a comprehensive survey for up-to-date ML-based automatic depression assessment (ADA) approaches building on various internal human brain behaviours recorded by EEG and MRIs, as well as external human expressive speech, facial, body and language behaviours. While early ADA surveys [71, 25, 72, 73] fail to include up-to-date advanced ADA approaches, recent surveys suffer from the following limitations: (i) reviewing only speech-based ADA solutions [74]; (ii) reviewing audio and language-based solutions without providing comprehensive reviews for visual ADA approaches [75]; and (iii) summing ADA approaches that analyse social media cues, ignoring audio, video, MRI and EEG-based ADA solutions [76]. More importantly, none of these surveys provided comprehensive psychological/physiological/medical evidence linking depression to various human internal and external behaviours. Our survey starts with reviewing the biological, physiological and psychological evidence linking depression with a wide variety of human internal signals and external behavioural cues (Sec. 2.2), which provide a solid theoretical basis for developing human internal and external cue-based ADA approaches (surveyed in Sec. 4). In addition, publicly available ADA datasets and competitions are discussed in Sec. 5. Finally, Sec. 6 discusses the challenges that are faced by existing ADA approaches and highlights opportunities for future research.
2 Background
@@
2.1 Health and social impacts of depression
Figure 1: The negative impacts of depression.
As summarised by Fig. 1, depression affects both human health and their social functioning.
(1) Health impacts: Depression frequently causes damages on human brain regions [77], particularly in the hippocampus, cortical regions, frontallobes, and thalamic [78]. It also affects human immune systems [79], leading to physiological illness such as coronary heart diseases [5] and cytokines-related illness [80]. A systematic survey [81] concludes that depressed chronically ill patients (e.g., diabetes, pulmonary disease and arthritis) suffer from significantly higher number of medical symptoms compared to non-depressed patients. Depression is also a high risk factor leading to peripheral vascular disease [82], which is also linked to chronic fatigue [7], insomnia [9], atherosclerosis [83], and Parkinson’s disease [84]. More seriously, depression is a potential risk factor for the initiation of cancers [85]. The psychological intervention treating depression can prolong survival of cancer patients with different types of cancers [86] while the meta-analysis conducted in [87] concluded that depression is a predictor of mortality in cancer patients.
@@ 2.1 Health and social impacts of depression
(2) Societal impacts: Depression also leads to serious social problems. More than half of the suicide cases in Europe are linked to depression-related mental disorders [88], with another study [89] revealed similar findings, i.e., the combined lifetime prevalence of depression and suicide attempts is 31%. Li et al. [90] surveyed 24 studies including 954,882 patients with depression, suggesting that depressed patients who experienced adverse environmental factors were more prone to suicidal ideation and attempting suicide. Depression also typically leads to poor interpersonal relationships [91, 92]. The inadequate interpersonal relationships, in turn, signify diminished social support and contribute to suicide-related behaviours [93]. Besides, addictive disorders are also frequently presented within depressed adolescents [94]. Moreover, depressed individuals frequently suffer from substantial financial costs in the form of time away from work and a rise in healthcare expenses [1], who are more susceptible to reduced workplace efficiency [6] diminished earnings/job loss, homeless and increased substance abuse and poverty [92].
2.2 Depression-related human behaviours
Previous studies found that depressed and non-depressed individuals frequently show different internal and external behaviours (illustrated in Fig. 2), which form the theoretical basis for developing existing ADA systems.
2.2.1 Brain cues
@@
Depressed individuals typically experience abnormal brain processes. This is evidenced by brain regions implicated in the pathophysiology of depression [95]. In particular, magnetic resonance imaging (MRI) have been widely employed to explore brain activities related to depression, revealing that depression would seriously impair several brain regions, including abnormal gray matter changes of dorsolateral prefrontal cortex (DLPFC) [77], amygdala [3] and cerebellum [96], as well as damage symptoms of hippocampus and thalamus (significantly smaller volumes [2, 78]). Some EEG studies found that depressed individuals usually exhibit hemispheric asymmetry in their brain signals compared to non-depressed individuals [97], as they usually have more alpha activity at the left side of their brains [98]. Besides, Li et al. [99] apply microstate analysis on EEG data of depressed and healthy individuals, suggesting that depression leads to increase in microstate C and decrease in microstate D, which further convinces depression’s negative influences on brain alpha band.